# Integrating Sequential and Relational Modeling for User Events: Datasets and Prediction Tasks

### v1

@@ 4.2 Prediction Tasks
Relational event prediction tasks. The corresponding tasks for pres-brightkite and pres-gowalla involve friend recommendation. We construct the training data by randomly splitting all relational events into 70% training, 10% validation, and 20% test sets. We also generate negative samples for the validation and test sets. Following Gastinger et al. [25], we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set. For the pres-github dataset, the relational event prediction task is defined as collaboration prediction, which involves predicting which users collaborate with a given user. The train, validation, and test splits follow the same procedure as in pres-brightkite, including the sampling method. However, due to the large size of the dataset, we adopt a 1-vs-300 negative sampling scheme.
@@
Personal event prediction tasks. The task for pres-brightkite and pres-gowalla is to predict the likelihood of a user checking in at a given geohash location in the future. We split each user’s personal events by taking the last 20% for test, the previous 10% for validation, and the rest for training. We also cap the number of events in the test and validation sets to at most 20 and 10 per user, respectively. Since personal events are more frequent than relational ones, we adopt a 1-vs-500 negative sampling scheme. As geohash strings encode hierarchical spatial information (e.g., earlier characters represent broader regions), we apply stratified hierarchical sampling. Specifically, negatives are stratified by shared geohash prefixes, from matching the first five characters to none, ensuring a mix of nearby and distant locations.
For the pres-amazon datasets, the task is to predict future products a user will review and the corresponding ratings, as denoted in their personal event data. We adopt the same train/val/test split strategy as in pres-brightkite, along with a 1-vs-500 negative sampling scheme. Negative samples for each personal event (e.g., B001OE3F08:3) are drawn from three sources: (1) the same product with different ratings (e.g., B001OE3F08:5); (2) other personal events not in the user’s training data; and (3) samples from the second set with randomly perturbed ratings. In the pres-github dataset, the number of unique events in the personal event set is only 24, corresponding to the list of possible GitHub activities. Thus, the task construction used in the previous datasets is not applicable to pres-github. We decided to omit this dataset from the set of datasets used for creating personal event prediction tasks.
Full event sequence. In addition to the datasets containing prediction tasks described above, we also publish a version of each dataset that includes all personal and relational events for all users, without any assigned tasks, train/val/test splits, or pre-specified negative samples. This is intended to facilitate future works that may wish to generate other prediction tasks not covered in this paper.
@@ 5.1 Relational event prediction tasks
    In the first set, we use only relational event data. We construct a user graph where edges represent relational events between two users, ignoring timestamp information. We then run static graph methods, GCN [45], GAT [46], and Graph Transformer (TConv) [47] on this graph.
@@
    In the second set, we use a sequence model, BERT [48], to encode each user’s last 100 personal events from the training set. The resulting user embedding is added as input to the GCN, GAT, and GT models from the first set, denoted as GCN+S, GAT+S, and TConv+S, respectively.
@@
    Lastly, based on the graph containing user and personal event nodes from the third set, we add timestamp information to construct a temporal graph. For datasets that lack timestamps for relational events, we inject these events randomly into the sequence of personal events. We then run temporal graph models, TGN [16], DyRep [19], and TNCN [20] on this graph.
@@
TGN, DyRep, and TNCN: Out of GPU Memory
For performance evaluation, following prior benchmarks [24, 25, 37], we use ranking-based metrics: Mean Reciprocal Rank (MRR) and Hits@k, evaluated at various $k$ depending on the number of negative samples. Each baseline is run five times with different random seeds, and we report the mean and standard deviation.
@@
  • In nearly all cases, converting personal events into nodes and adding them to the relational event graph is less effective than modeling personal events with a sequence model and using the user’s sequence embedding as an additional node features to the relational event graph. The performance of the “+S” versions of static graph models is lower than that of the corresponding “_RP” versions across nearly all datasets and metrics, with a few exceptions on the amazon datasets for Hits@k metrics with lower $k$.
@@ 5.2 Personal event prediction tasks
    The first model is a sequential model that uses only personal event data. We use a BERT architecture with a prediction head to compute the likelihood of a user having a particular personal event in the future. For each user, we use the last 100 personal events in the training set to predict the likelihood of future events.
@@
    In the second set, we use node2vec [49] to learn the graph structure of relational events and generate a graph embedding for each user. We then incorporate the embedding into the BERT sequence model. We evaluate two versions of the model: (a) incorporating the graph embedding post transformer module and before the prediction head (BERT-n2v-p), and (b) using the embedding as a special input token to the transformer module (BERT-n2v-i).
@@
Similar to the sequence embedding used in relational event prediction tasks, we apply split tokenization for the BERT model in personal event prediction to allow more flexibility in modeling events. We use the same tokenization scheme for each dataset as described earlier. For evaluation, we report MRR and Hits@k at various values of $k$. Each baseline is run five times with different random seeds, and we report the mean and standard deviation.
@@
6 Conclusions
@@ Appendix A Dataset Documentation
  • pres-amazon-clothing and pres-amazon-electronics. These datasets contain Amazon product reviews and ratings in their respective categories. Both are based on Amazon review data collected by McAuley et al. [44] and hosted at: https://cseweb.ucsd.edu/˜jmcauley/datasets/amazon/links.html. The Amazon review content is licensed under the Amazon license:
@@
  • pres-gowalla. This dataset contains user activity on the (now defunct) social network Gowalla. It was originally collected by Cho et al. [40] using the platform’s public API and published in the SNAP Dataset Repository [41] (https://snap.stanford.edu/data/loc-Gowalla.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
  • pres-brightkite. This dataset contains user activity on the (also now defunct) social network Brightkite. It was also originally collected by Cho et al. [40] using the platform’s public API and published in the SNAP Dataset Repository [41] (https://snap.stanford.edu/data/loc-brightkite.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
@@ Appendix C Dataset Construction
To generate the relational event prediction task, we perform a random split of 70%/10%/20% on the relational event dataframe to generate the train/eval/test datasets, and store them into CSV files. Following Gastinger et al. [25], we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set.
@@ Appendix D Additional Related Works
Temporal point processes (TPPs) provide a powerful tool for modeling discrete events as stochastic processes in continuous time. Classical models like the Poison process and the Hawkes process [50] allow for explicit modeling of event dynamics, including self-excitation and mutual inhibition, through parameterized intensity functions that govern the likelihood of future events. Recent work has introduced neural extensions such as Neural Hawkes Process [26] [27] and the Self-Attentive Hawkes Process [28], which integrate RNNs or attention mechanisms into the intensity function. These models are particularly well suited for fine-grained timestamp prediction and have found applications in finance, healthcare, and user behavior modeling. However, TPPs typically assume a simple structure over events and focus solely on the temporal dimension, making them less suitable for capturing structured dependencies across users or networks.
@@
A closely related application domain is sequential recommendation, where the goal is to predict the next item a user will interact with based on their historical behavior. Early methods employed Markov chains or matrix factorization over time-slided data [51], while more recent models use deep sequence encoders such as GRU4Rec [32], SASRec [33], and BERT4Rec [13], which apply Transformer-based architectures to item sequences. These models have shown strong performance by capturing user preferences over time. However, they typically model each user independently and do not account for interactions among users.
@@
In parallel, there has been significant progress in graph-based modeling of user interaction, especially through Graph Neural Networks (GNN). While static graphs lack explicit timestamps and do not capture the order of interactions, they can still represent temporal data through careful graph construction, such as building graphs over specific time windows or pruning outdated edges [52]. This setup allows for framing personal and relational event prediction tasks as link prediction (e.g. forecasting probability of user-item interactions) or node classification (e.g. fraud detection). Early GNNs like GCN [45] introduced neighborhood aggregation but were limited by their transductive nature and the requirement of full graph knowledge during training. GraphSAGE [53] addressed these issues by introducing inductive learning via stochastic sampling, while PinSAGE [54] scaled GNNs to billions of nodes via relaxing memory constraints. GAT [46] incorporated attention mechanisms and later HGT [55] extends GAT to heterogeneous graphs. These developments have established GNNs as a powerful tools for relational modeling [10].
@@
As described in Gastinger et al. [25], temporal graph methods fall into two broad categories: discrete-time and continuous-time. Discrete-time methods exist for both homogeneous [56] and heterogeneous datasets [57, 58, 59]. Continuous-time methods arguably preserve more information and can be converted into discrete graphs, but the reverse is not possible [17]. TGN [16] introduces a general framework for modeling continuous-time dynamic graphs, categorizing DyRep [19] as a special case, followed by several other models such as DyGFormer [22], NAT [21], TNCN [20], and CTAN [23]. Other temporal graph models such as HTGN-BTW [60] and STHN [61] propose different extensions of TGN to handle heterogeneous data. Beyond the standard temporal graphs, several methods have also been proposed for modeling temporal knowledge graphs [62, 63, 64].
@@
Several benchmark efforts have been proposed across related areas. The temporal graph benchmarks include the TGB [24], its heterogeneous and knowledge graph extension (TGB 2.0) [25], and TGB-Seq [37], which include a more complex sequence of edge dynamics. For static graph learning, OGB [38] and its large-scale extension OGB-LSC [39] provide widely used benchmarks. In the recommendation domain, large-scale user-item interaction datasets have been released through benchmarks such as MIND [65], TenRec [66], NineRec [67], and BARS [68]. For event sequence modeling and temporal point processes (TPP), recent benchmarks include EBES [69], EasyTPP [70], and HOTPP [71].
The closest dataset or benchmark work from our paper is the temporal graph benchmark (TGB) by Huang et al. [24]. It contains several datasets that model user behaviors. These datasets can be roughly divided into two categories: (1) user-to-user interaction datasets, such as tgbl-coin, tgbl-comment, and tgbn-trade; (2) user-to-item bipartite interaction datasets, such as tgbl-wiki, tgbl-review, tgbn-genre, tgbn-reddit, and tgbn-token. The first category of user-to-user interaction datasets are similar to the relational event part of our datasets; however our datasets also contain personal event sequences, in addition to relational events, which are not present in the TGB dataset.
@@
Several studies have been conducted on different settings on temporal and structural dynamics. Some models focus on modeling graph and time series data using spatio-temporal graph [72, 73, 74, 75]. Other model combine graph model output and sequence model output in various application areas [76, 77, 78]. A recent study focus on tokenizing graph and applying transformers or state space models (SSMs) for graph learning [79, 80, 81, 82, 83, 84, 85]. Another studies works on incorporating knowledge graph into language model. [86, 87, 88], as well as performing graph-augmented language generation [89, 90].

### v2

@@ 4.2 Prediction Tasks
Relational event prediction tasks.
The corresponding tasks for pres-brightkite and pres-gowalla involve friend recommendation. We construct the training data by randomly splitting all relational events into 70% training, 10% validation, and 20% test sets. We also generate negative samples for the validation and test sets. Following gastinger2024tgb, we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set.
For the pres-github dataset, the relational event prediction task is defined as collaboration prediction, which involves predicting which users collaborate with a given user. The train, validation, and test splits follow the same procedure as in pres-brightkite, including the sampling method. However, due to the large size of the dataset, we adopt a 1-vs-300 negative sampling scheme.
@@
Personal event prediction tasks.
The task for pres-brightkite and pres-gowalla is to predict the likelihood of a user checking in at a given geohash location in the future. We split each user’s personal events by taking the last 20% for test, the previous 10% for validation, and the rest for training. We also cap the number of events in the test and validation sets to at most 20 and 10 per user, respectively. Since personal events are more frequent than relational ones, we adopt a 1-vs-500 negative sampling scheme. As geohash strings encode hierarchical spatial information (e.g., earlier characters represent broader regions), we apply stratified hierarchical sampling. Specifically, negatives are stratified by shared geohash prefixes, from matching the first five characters to none, ensuring a mix of nearby and distant locations.
For the pres-amazon-clothing and pres-amazon-electronics datasets, the task is to predict future products a user will review and the corresponding ratings, as denoted in their personal event data. We adopt the same train/val/test split strategy as in pres-brightkite, along with a 1-vs-500 negative sampling scheme. Negative samples for each personal event (e.g., B001OE3F08:3) are drawn from three sources: (1) the same product with different ratings (e.g., B001OE3F08:1, B001OE3F08:5); (2) other personal events not in the user’s training data; and (3) samples from the second set with randomly perturbed ratings.
In the pres-github dataset, the number of unique events in the personal event set is only 24, corresponding to the list of possible GitHub activities. Thus, the task construction used in the previous datasets is not applicable to pres-github. We decided to omit this dataset from the set of datasets used for creating personal event prediction tasks.
Full event sequence.
In addition to the datasets containing prediction tasks described above, we also publish a version of each dataset that includes all personal and relational events for all users, without any assigned tasks, train/val/test splits, or pre-specified negative samples. This is intended to facilitate future works that may wish to generate other prediction tasks not covered in this paper.
@@ 5.1 Relational event prediction tasks
    In the first set, we use only relational event data. We construct a user graph where edges represent relational events between two users, ignoring timestamp information. We then run static graph methods, GCN [kipf2017semi], GAT [velivckovic2018graph], and Graph Transformer (TConv) [shi2021masked] on this graph.
@@
    In the second set, we use a biderectional transformer sequence model (BERT) [devlin2019bert], to encode each user’s last 100 personal events from the training set. The resulting user embedding is added as input to the GCN, GAT, and TConv models from the first set, denoted as GCN+S, GAT+S, and TConv+S, respectively.
@@
    Lastly, based on the graph containing user and personal event nodes from the third set, we add timestamp information to construct a temporal graph. For datasets that lack timestamps for relational events, we inject these events randomly into the sequence of personal events. We then run temporal graph models, TGN [tgn_icml_grl2020], DyRep [trivedi2019dyrep], and TNCN [zhang2406efficient] on this graph.
@@
For performance evaluation, following prior benchmarks [huang2023temporal, gastinger2024tgb, yi2025tgb], we use ranking-based metrics: Mean Reciprocal Rank (MRR) and Hits@k, evaluated at various $k$ depending on the number of negative samples. Each baseline is run five times with different random seeds, and we report the mean and standard deviation of the results.
@@
TGN Out of GPU Memory
DyRep Out of GPU Memory
TNCN Out of GPU Memory
@@
  • In nearly all cases, converting personal events into nodes and adding them to the relational event graph is less effective than modeling personal events with a sequence model and using the user’s sequence embedding as an additional node features to the relational event graph. The performance of the “_RP” versions of static graph models is lower than that of the corresponding “+S” versions across nearly all datasets and metrics, with a few exceptions on the amazon datasets for Hits@k metrics with lower $k$.
@@ 5.2 Personal event prediction tasks
    The first model is a sequential model that uses only personal event data. We use a bidirectional transformer architecture (BERT) with a prediction head to compute the likelihood of a user having a particular personal event in the future. For each user, we use the last 100 personal events in the training set to predict the likelihood of future events.
@@
    In the second set, we use node2vec [grover2016node2vec] to learn the graph structure of relational events and generate a graph embedding for each user. We then incorporate the embedding into the BERT sequence model. We evaluate two versions of the model: (a) incorporating the graph embedding post transformer module and before the prediction head (BERT-n2v-p), and (b) using the embedding as a special input token to the transformer module (BERT-n2v-i).
@@
Similar to the sequence embedding used in relational event prediction tasks, we apply split tokenization for the BERT model in personal event prediction to allow more flexibility in modeling events. We use the same tokenization scheme for each dataset as described earlier. For evaluation, we report MRR and Hits@k at various values of $k$. Each baseline is run five times with different random seeds, and we report the mean and standard deviation of the results.
@@
5.3 Hypothesis on native end-to-end model architectures for PRES datasets
In the experiment results described above, we have demonstrated that current models leave notable room for performance improvements in both personal and relational prediction tasks. We hypothesize that a significant performance improvement can be achieved by end-to-end models trained simultaneously on both personal and relational event data. There are many ways to design such models; below are some examples:
  1. 1.
    A possible strategy is to start with a user-level Transformer architecture that takes the user event sequence as input, combined with a message-passing architecture that enables information transfer between two users connected by a relational event and propagates that information across the neighbors via a Graph Neural Network (GNN)-like architecture. This design capitalizes on the strengths of both the Transformer and GNN architectures. Transformer architectures have been shown to be the best model for handling sequential information, whereas GNN architectures excel in transferring information across connected users.
  2. 2.
    Another possible architecture candidate is to also start from the Transformer architecture for the sequence, where each sample represents a user event sequence, and allow the attention mechanism to propagate toward other user sequences that are connected via a relational event, in addition to the attention mechanism inside each user/sequence.
There could also be multiple other promising architectural paradigms for developing a fully end-to-end model. We leave the exploration of these alternative approaches for future work.
6 Conclusions and Limitations
@@
A key challenge in this work is in the dataset curation process, as many public datasets have already been collapsed into either graph-only or sequence-only formats, often discarding personal or relational events in the process. While we were able to gather and unify a set of datasets that include both event types, they may not fully capture the diversity and complexity of user event modeling across domains. Another limitation is that our current formulation does not support event-level or user-level features, presenting an opportunity for future work to extend the framework toward feature-aware modeling.
@@ Appendix A Dataset Documentation
  • pres-amazon-clothing and pres-amazon-electronics. These datasets contain Amazon product reviews and ratings in their respective categories. Both are based on Amazon review data collected by mcauley2015image and hosted at: https://cseweb.ucsd.edu/˜jmcauley/datasets/amazon/links.html. The Amazon review content is licensed under the Amazon license:
@@
  • pres-gowalla. This dataset contains user activity on the (now defunct) social network Gowalla. It was originally collected by cho2011friendship using the platform’s public API and published in the SNAP Dataset Repository [snapnets] (https://snap.stanford.edu/data/loc-Gowalla.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
  • pres-brightkite. This dataset contains user activity on the (also now defunct) social network Brightkite. It was also originally collected by cho2011friendship using the platform’s public API and published in the SNAP Dataset Repository [snapnets] (https://snap.stanford.edu/data/loc-brightkite.html). The curator confirmed that SNAP datasets are free to use, but no specific license information is available.
@@ Appendix C Dataset Construction
To generate the relational event prediction task, we perform a random split of 70%/10%/20% on the relational event dataframe to generate the train/eval/test datasets, and store them into CSV files. Following gastinger2024tgb, we adopt a 1-vs-1000 negative sampling scheme, in which 1,000 negative events are sampled for each relational event in the prediction set. Negative samples are drawn via uniform random sampling of users, excluding those who already have relational events with the target user in the training set.
@@ Appendix D Additional Related Works
Temporal point processes (TPPs) provide a powerful tool for modeling discrete events as stochastic processes in continuous time. Classical models like the Poison process and the Hawkes process [hawkes1971spectra] allow for explicit modeling of event dynamics, including self-excitation and mutual inhibition, through parameterized intensity functions that govern the likelihood of future events. Recent work has introduced neural extensions such as Neural Hawkes Process [mei2017neural] [zuo2020transformer] and the Self-Attentive Hawkes Process [zhang2020self], which integrate RNNs or attention mechanisms into the intensity function. These models are particularly well suited for fine-grained timestamp prediction and have found applications in finance, healthcare, and user behavior modeling. However, TPPs typically assume a simple structure over events and focus solely on the temporal dimension, making them less suitable for capturing structured dependencies across users or networks.
@@
A closely related application domain is sequential recommendation, where the goal is to predict the next item a user will interact with based on their historical behavior. Early methods employed Markov chains or matrix factorization over time-slided data [rendle2010factorizing], while more recent models use deep sequence encoders such as GRU4Rec [hidasi2015session], SASRec [kang2018self], and BERT4Rec [sun2019bert4rec], which apply Transformer-based architectures to item sequences. These models have shown strong performance by capturing user preferences over time. However, they typically model each user independently and do not account for interactions among users.
@@
In parallel, there has been significant progress in graph-based modeling of user interaction, especially through Graph Neural Networks (GNN). While static graphs lack explicit timestamps and do not capture the order of interactions, they can still represent temporal data through careful graph construction, such as building graphs over specific time windows or pruning outdated edges [zhao2025gigl]. This setup allows for framing personal and relational event prediction tasks as link prediction (e.g. forecasting probability of user-item interactions) or node classification (e.g. fraud detection). Early GNNs like GCN [kipf2017semi] introduced neighborhood aggregation but were limited by their transductive nature and the requirement of full graph knowledge during training. GraphSAGE [hamilton2017inductive] addressed these issues by introducing inductive learning via stochastic sampling, while PinSAGE [ying2018graph] scaled GNNs to billions of nodes via relaxing memory constraints. GAT [velivckovic2018graph] incorporated attention mechanisms and later HGT [hu2020heterogeneous] extends GAT to heterogeneous graphs. These developments have established GNNs as a powerful tools for relational modeling [gao2023survey].
@@
As described in gastinger2024tgb, temporal graph methods fall into two broad categories: discrete-time and continuous-time. Discrete-time methods exist for both homogeneous [han2020graph] and heterogeneous datasets [yin2019dhne, xue2021modeling, bian2019network]. Continuous-time methods arguably preserve more information and can be converted into discrete graphs, but the reverse is not possible [kazemi2020representation]. TGN [tgn_icml_grl2020] introduces a general framework for modeling continuous-time dynamic graphs, categorizing DyRep [trivedi2019dyrep] as a special case, followed by several other models such as DyGFormer [yu2023towards], NAT [luo2022neighborhood], TNCN [zhang2406efficient], and CTAN [gravina2024long]. Other temporal graph models such as HTGN-BTW [yue2022htgn] and STHN [li2023simplifying] propose different extensions of TGN to handle heterogeneous data. Beyond the standard temporal graphs, several methods have also been proposed for modeling temporal knowledge graphs [li2021temporal, jin2019recurrent, li2022complex].
@@
Several benchmark efforts have been proposed across related areas. The temporal graph benchmarks include the TGB [huang2023temporal], its heterogeneous and knowledge graph extension (TGB 2.0) [gastinger2024tgb], and TGB-Seq [yi2025tgb], which include a more complex sequence of edge dynamics. For static graph learning, OGB [hu2020open] and its large-scale extension OGB-LSC [hu2021ogb] provide widely used benchmarks. In the recommendation domain, large-scale user-item interaction datasets have been released through benchmarks such as MIND [wu2020mind], TenRec [yuan2022tenrec], NineRec [zhang2024ninerec], and BARS [zhu2022bars]. For event sequence modeling and temporal point processes (TPP), recent benchmarks include EBES [osin2024ebes], EasyTPP [xue2023easytpp], and HOTPP [karpukhin2024hotpp].
The closest dataset or benchmark work from our paper is the temporal graph benchmark (TGB) by huang2023temporal. It contains several datasets that model user behaviors. These datasets can be roughly divided into two categories: (1) user-to-user interaction datasets, such as tgbl-coin, tgbl-comment, and tgbn-trade; (2) user-to-item bipartite interaction datasets, such as tgbl-wiki, tgbl-review, tgbn-genre, tgbn-reddit, and tgbn-token. The first category of user-to-user interaction datasets are similar to the relational event part of our datasets; however our datasets also contain personal event sequences, in addition to relational events, which are not present in the TGB dataset.
@@
Several studies have been conducted on different settings on temporal and structural dynamics. Some models focus on modeling graph and time series data using spatio-temporal graph [yu2017spatio, yu2020spatio, bai2020adaptive, wu2020connecting]. Other model combine graph model output and sequence model output in various application areas [hai2024study, shi2019two, mao2021molecular]. A recent study focus on tokenizing graph and applying transformers or state space models (SSMs) for graph learning [zhang2001graph, wang2024graph, behrouz2024graph, huang2024can, ding2024dygmamba, li2024dyg, behrouz2024best]. Another studies works on incorporating knowledge graph into language model. [liu2020k, yasunaga2021qa, huang2020knowledge], as well as performing graph-augmented language generation [peng2024graph, zhu2025graph].